OpenStudy (iwanttogotostanford):

Help Please:

OpenStudy (iwanttogotostanford):

@imqwerty @sweetburger

OpenStudy (luffingsails):

cot (x - pi/2) = cos( x-pi/2) / sin(x - pi/2) cos(a-b) = cos(a)cos(b) +sin(a)sin(b) cos(x-pi/2) = cos(x) cos(pi/2) + sin(x) sin(pi/2) cos(x-pi/2) = cos(x) (0) + sin(x) (1) = sin (x) sin(a-b) = sin(a)cos(b) -cos(a)sin(b) sin(x-pi/2) = sin(x) cos(pi/2) - cos(x) sin(pi/2) sin(x-pi/2) = sin(x) (0) = -cos(x) (1) cos(x-pi/2) / sin(x-pi/2) = sin(x) / [-cos(x)] = -tan(x)

OpenStudy (luffingsails):

There is an easier way but that is probably the one that you are looking for.

OpenStudy (faiqraees):

Yes, are you aware of the double angle formula?\[\large\rm tan(\alpha+\beta)= \frac{tan(\alpha)+tan(\beta )}{1-tan(\alpha)tan(\beta )}\]

OpenStudy (iwanttogotostanford):

thank you both! yes, I'm aware